OpenStudy (anonymous):

Lorenzo asked his friends, “What’s your favorite instrument?” He recorded the results in a frequency table. Calculate relative frequency to determine the answer. Based on data in the table, which statement about relative frequency is true? Favorite Instrument Instrument guitar trumpet drums saxophone piano Frequency 12 3 10 3 2 Relative frequency A. Bass is 30%. B. Guitar is 12%. C. Saxophone is 10%. D. Piano is 2%.

OpenStudy (anonymous):

@iGreen @Godlovesme

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Awolflover1

Awolflover1 (awolflover1):

@acxbox22

OpenStudy (acxbox22):

what do you think?

OpenStudy (anonymous):

I think it's C

OpenStudy (acxbox22):

ya C is right once you add all of them up we get 30 3/30=1/10=10%

OpenStudy (anonymous):

Okie thx :)